OpenStudy (anonymous):

At what points do the given equations intersect: r1=<1,1,0>+t<1,-1,2> r2=<2,0,2>+s<-1,1,0>

OpenStudy (anonymous):

I already found that s=0 and t=1 when they intersect, but I don't know how to find the exact point from that.

jimthompson5910 (jim_thompson5910):

r1=<1,1,0>+t<1,-1,2> r1=<1,1,0>+1<1,-1,2> r1=<1,1,0>+<1*1,1*(-1),1*2> r1=<1,1,0>+<1,-1,2> r1=<1+1,1-1,0+2> r1=<2,0,2> ------------------------------------------------------- OR r2=<2,0,2>+s<-1,1,0> r2=<2,0,2>+0<-1,1,0> r2=<2,0,2>+<0*(-1),0*1,0*0> r2=<2,0,2>+<0,0,0> r2=<2+0,0+0,2+0> r2 = <2,0,2> Either way, we get <2,0,2>

OpenStudy (anonymous):

Thank you!

jimthompson5910 (jim_thompson5910):

np